OpenStudy (anonymous):

the equation y=1777x+27.153 can be used to predict the number y of gun death in the united states x years after 2000that x=0 correspound to 2000,x=3 correspound to 2006 and so on .predict the number of gun death in 2006 and 2007 in what year will the number of guns deaths be 12,937

OpenStudy (anonymous):

the predict of gun death in 2006 will be ?

OpenStudy (anonymous):

the predicated of gun death 2007 will be ?

OpenStudy (anonymous):

well, the problem itself sates that x=3 corresponds to 2006, and the equation is used to predict the number of deaths in the US x years after 2000. This means that all we have to to is simply plug in x=3 into out equation like so: f(3)=1777(3)+27.153

OpenStudy (anonymous):

you have to evaluate that to get your answer

OpenStudy (anonymous):

The predicate a number of gun death will be 12.937in the year

OpenStudy (anonymous):

the same can be done to predict the number of deaths in 2007, now here notice that x=3 corresponded to 2006, what does that tell you? from 2000 to 2006, 6 years has passed, but x=3 is assigned to the year 2006. That means that we are going by increments of 2, you see 6/3 =2. So in terms of x, 1=2years. Now since x=3 corresponds to 2006, then naturally 2007 must correspond to x=4. Like we did above we plug in x=4 into the equation: f(4)=1777(4)+27.153 Evaluate that for your answer(number of deaths in 2007)

OpenStudy (anonymous):

now if we want to predict the year that the number dead will be 12,937, we simply have to set our equation equal to this number and solve for x(the year) like so: 12,937=1777x+27.153, now i am confident that you can solve for x. This will give you the approximat year where this number of deaths occured

OpenStudy (anonymous):

so it would 6.7 2006

OpenStudy (anonymous):

for which one?

OpenStudy (anonymous):

for 2006 would b 6 )and for 2007 would b 7

OpenStudy (anonymous):

for 2006, the number of projected deaths would be about 5358.153

OpenStudy (anonymous):

I have a correction for the one about 2007, it would be x=3.5, since for every 1(interms of x) it is 2years. And since x=3 corresponded to 2006, 6 years after 2000. Then x=3.5 corresponds to 2007. So we have instead f(3.5)=1777(3.5)+27.153

OpenStudy (anonymous):

So, the projected number of deaths in 2007 would be : 6246.653

OpenStudy (anonymous):

now, is it is the case that we are not getting the same answer, then you need to check your problem again to make sure that you copied it correctly